It takes a computer 8 days to print all of the letters for a sweepstakes. A new computer is purchased that can do the same job in 5 days. How long would it take both computers working on it.
Answer by checkley77(12844)

8*5/(8+5)=40/13=3.077 hours to print the letters working together.